### v1

# Towards grid-aware multi-period flexibility aggregation – A constrained zonotope approach

Abstract
Aggregation schemes provide a means to reduce the computational complexity of power system operation by reducing the number of devices that are considered individually. This can be achieved with tools of computational geometry, where the feasible set is projected onto the decision variables of the point of interconnection. Set projection is computationally expensive, especially in the context of multi-period power system operation. Hence this calls for efficiency improvements via structure exploitation of certain set representations, such as constrained zonotopes. This paper proposes these benefits for efficient flexibility aggregation. We evaluate the performance of the proposed method on a 15-bus distribution grid with time-dependent elements for up to 96 timesteps. The results suggest that the presented method significantly improves computation times.
I Introduction
The increasing integration of Distributed Energy Resources (DERs) into modern power systems introduces new operational challenges for both Transmission System Operators (TSOs) and Distribution System Operators (DSOs). From the perspective of the TSO, higher volatility and the spatial distribution of generation increase the risk of voltage violations, line congestion, and inverted power flows. Moreover, most flexible devices are located in distribution grids added at the distribution level. The availability of DERs and storage solutions in the distribution grid offers new opportunities for flexible grid operation, which in turn creates new coordination requirements between the TSO and the DSOs to ensure reliable and efficient system performance.
These interactions can be modeled through a centralized Optimal Power Flow (OPF) formulation, in which all system data—such as DER capacities, grid parameters, and asset locations—are collected by a central operator that solves a single global optimization problem. However, to preserve data privacy, reduce computational complexity and to simplify information exchange, such detailed information from the distribution grids should not be directly shared with the TSO.
To address these challenges, hierarchical optimization has been widely adopted in the literature. In the first stage of the approach, flexibility aggregation, the Feasible Operation Region (FOR) is computed. The FOR describes the values of the shared variables, usually active and reactive power, at the interconnection point that can be achieved without compromising the feasibility of the optimization problem from the DSO’s perspective. A survey of different methods can be found in [1]. The TSO can then use this information to determine the optimal utilization of DER flexibility without requiring direct data from the distributed grid assets.
The standard FOR estimation methods rely on sampling the region and computing the convex hull of these points. The samples can be randomly obtained [2] or calculated in a more structured manner, utilizing knowledge of the grid and constraints to solve optimization problems [3, 4, 5]. The methods above consider assumptions that can be quite limiting. They assume a unique interconnection point between DSO and TSO, the voltage at this point to be constant, and no time-dependent elements, e.g., storage systems.
There is limited research on the flexibility aggregation in distribution grids that considers time-dependent elements. A survey of various methods for aggregating batteries can be found in [6]. They offer computational tractability but neglect grid constraints or rely on data-driven formulations, which reduces their accuracy for grid-aware applications. Multi-period cost curve formulations [7] capture temporal coupling but limits the FOR to only portraying active power.
Recently, set projection has been used for the computation of the FOR [8, 9]. The method has been extended to FORs, which include time-dependent elements [10], but it requires solving multiple optimization problems. For affine grid models, such as LinDistFlow [11], the projection can be obtained in one step via tools of computational geometry. The downside of this technique is the computational burden of the set projection operation, which limits its scalability [12].
One line of research addresses this computational burden of polytope projection by applying zonotope approximations [13, 14]. However, such approximations are not always accurate, particularly when the underlying set is non-symmetric.
The present paper introduces a novel approach to compute time-coupled FORs by extending the non-iterative projection-based formulation [15] to multi-period settings. To overcome computational bottlenecks that normally hinder the scalability of set projections, we rely on Constrained Zonotopes ($\mathcal{CZ}$[16]. CZs enable non-symmetric set representation with efficient set projection properties. This results in significant reduction of computation times without compromising representation accuracy. Furthermore, most of the remaining computation can be carried out offline, enabling the practical implementation for the computation of FORs for a 15-bus distribution grid that include time-dependent elements such as storage units for up to 96 timesteps.
The remainder of this paper is structured as follows: Section II elaborates the problem statement and recaps the flexibility aggregation in the ADP approach. In Section III we present how $\mathcal{CZ}\mathrm{s}$ can be utilized for efficient flexibility aggregation over time, while Section IV highlights the potential of this method for dynamic problem settings. Section V compares and discusses numerical results for the presented methods. The paper concludes with Section VI.
II Problem Statement
We assume a tree-structured power system, divided into transmission and distribution grids, as shown in Figure 1. Time coupled system operation across multiple grid levels requires the solution of multi-stage Optimal Power Flow (OPF) problems. Next, we introduce one variant of this problem, which we later use as a basis for aggregation. We limit the presentation to one distribution grid in order to simplify presentation.
The first step in flexibility aggregation via set projection is to define the constraint set of the sub-problems, i.e. the distribution grids. We represent the distribution grid as a graph $G^{e}=(\mathcal{N}_{d},\mathcal{B}_{d})$, where $\mathcal{N}_{d}$ is the set of all buses and $\mathcal{B}_{d}\subseteq\mathcal{N}_{d}\times\mathcal{N}_{d}$ is the set of branches, i.e., lines and transformers.
We define the nodal active and reactive power $p_{m}(k)$ and $q_{m}(k)$ at each timestep $k$ in the considered interval $\mathcal{I}={1,\dots,N}$ as
$p_{m}(k)=\sum_{l\in\mathcal{N_{d}}}p_{m,l}(k),\qquad q_{m}=\sum_{l\in\mathcal{N}}q_{m,l}(k),$ (1)
and the net powers at each node
$p_{m}(k)$ $=p_{m}^{g}(k)-p_{m}^{d}(k)+p_{m}^{s}(k),$ (2a)
$q_{m}(k)$ $=q_{m}^{g}(k)-q_{m}^{d}(k),$ (2b)
where $p_{m}^{g}(k)$ and $q_{m}^{g}(k)$ are the active/reactive power generation and $p_{m}^{d}$ and $q_{m}^{d}(k)$ are the active/reactive power demand at node $m\in\mathcal{N}_{d}$. Furthermore, $p_{m}^{s}(k)$ is the battery’s active power. For nodes without a battery, $p_{m}^{s}=0$, while for nodes with a battery, $p_{m}^{s}$ is negative when charging and positive when discharging, i.e., providing the grid with power.
The active power generation for the flexible renewable generators must stay within the bounds
$0\leq p_{m}^{g}(k)\leq\bar{f}_{m}^{g}(k),\qquad m\in\mathcal{G}.$ (3)
where $\mathcal{G}\subseteq\mathcal{N}$ denotes the set of buses with controllable generators and $f_{m}^{g}(k)$ represents the maximal active power generation given by irradiation/wind forecast conditions at each timestep.
We limit the reactive power support from each renewable generator by constraining the maximum apparent power $\bar{s}_{m}(k)$ and the power factor limit $\alpha$ at each timestep with the affine constraints [17]
$p_{m}^{g}(k)\leq\bar{s}_{m}(k)\cos(\alpha),\quad-p_{m}(k)\leq\alpha q_{m}(k)\leq p_{m}(k).$ (4)
Next, we formulate the active/reactive power flow $p_{m,l}(k)$ and $q_{m,l}(k)$ on each line $(m,l)\in\mathcal{B}_{d}$ and the squared voltage $\nu_{m}$ at each bus $m\in\mathcal{N}_{d}$ according to the DistFlow formulation for radial grids [18]
$p_{m,l}(k)$ $=\sum_{j\in\mathcal{N}_{d}}p_{m,j}(k)+p_{m}(k)+r_{m,l}\,\ell_{m,l}(k),$ (5a)
$q_{m,l}(k)$ $=\sum_{j\in\mathcal{N}_{d}}q_{m,j}(k)+q_{m}(k)+x_{m,l}\ell_{m,l}(k),$ (5b)
$\nu_{m}(k)$ $=\nu_{l}(k)+2(r_{m,l}\,p_{m,l}(k)+x_{m,l}\,q_{m,l}(k))$
$-(r_{m,l}^{2}+x_{m,l}^{2})\ell_{m,l}(k),$ (5c)
where, $r_{m,l}$ and $x_{m,l}$ are, respectively, the line reactance and capacitance and $\ell_{m,l}(k)$ is the squared line current.
To obtain a polyhedral constraint set, we use the first-order Taylor series as linear approximation for the square current $\ell_{m,l}(k)$ around the operational point $t_{m,l}^{0}$ [19],
$\ell_{m,l}(k)\approx\ell_{m,l}(k)^{0}+J_{m,l}|_{\begin{subarray}{c}t_{m,l}^{0}(k)\end{subarray}}^{\intercal}\delta_{m,l}(k)$ (6)
where
$\delta_{m,l}(k)$ $=\begin{bmatrix}p_{m,l}(k)-p_{m,l}^{0}\\ q_{m,l}(k)-q_{m,l}^{0}\\ \nu_{l}(k)-\nu_{l}^{0}\end{bmatrix}$, (7)
$J_{m,l}|_{\begin{subarray}{c}t_{m,l}^{0}\end{subarray}}$ $=\begin{bmatrix}\frac{2p_{m,l}(k)}{\nu_{l}^{0}}\\ \frac{2q_{m,l}(k)}{\nu_{l}^{0}}\\ -\frac{(p_{m,l}^{0})^{2}+q_{m,l}^{0})^{2}}{(\nu_{l}^{0})^{2}}\end{bmatrix}$. (8)
Additionally, we limit the squared voltage and the squared current
$\underline{\nu}_{m}\leq\nu_{m}(k)\leq\bar{\nu}_{m},\quad\underline{\ell}_{m,l}\leq\ell_{m,l}(k)\leq\bar{\ell}_{m,l},$ (9)
Moreover, let $\mathcal{S}\subseteq\mathcal{N}$ denote the set of buses equipped with storage units. Each storage unit has a limited energy capacity $e_{m}^{s}(k)$, bounded by:
$\underline{e}^{s}_{m}\leq e^{s}_{m}(k)\leq\overline{e}^{s}_{m},\quad m\in\mathcal{S}$ (10)
with the maximum charging and discharging rate of
$\underline{p}^{s}_{m}\leq p_{m}^{s}(k)\leq\overline{p}^{s}_{m},\quad m\in\mathcal{S}.$ (11)
In this work we neglect the reactive power support of storage systems.
A simple discrete-time integrator dynamic model determines the State of Charge (SOC) of the storage unit. Assuming a fixed timestep size $\Delta k$ and neglecting losses, the SOC dynamic can be described as
$e_{m}(k+1)=e_{m}(k)-\Delta k\cdot p_{m}^{s}(k),\quad m\in\mathcal{S}.$ (12)
The initial condition of the battery is given by
$e_{m}(1)=e_{m}^{0},\quad m\in\mathcal{S}.$ (13)
We define the resulting constraint set as
[…]
$\forall k\in\mathcal{I},\,\text{and }\forall(m,l)\in\mathcal{B}_{m}\big\}.$
The vector of coupling variables with the TSO is defined as $z_{i}^{\top}=[[p_{i,j}(k),q_{i,j}(k)]_{i,j\in\mathcal{B}_{d},k\in\mathcal{I}},[\nu_{i}(k)]_{i\in\mathcal{N}_{d}^{c},k\in\mathcal{I}}]$. These variables appear in both the TSO and DSO optimization problems, representing the physical coupling between the two grids. Here, $\mathcal{N}_{d}^{c}$ Denotes the set of nodes that serve as coupling points between the transmission and distribution grids.
Figure 1: Tree-structured optimization of power system operation adapted from [15].
The resulting constraint set is a convex and bounded polyhedron, i.e. a polytope. The set projection of $\mathcal{X}\subseteq\mathcal{Y}\times\mathcal{Z}$ onto $\mathcal{Z}$ is defined as in [20]
$\operatorname{proj}_{\mathcal{Z}}(\mathcal{X})\doteq\{z\in\mathcal{Z}\ |\ \exists y\in\mathcal{Y}\text{ with }(z,y)\in\mathcal{X}\}\subseteq\mathcal{Z},$ (14)
which gives us the possible values for the coupling variables that can be achieved without violating any of the constraints of the original set, i.e., we can compute the FOR $\mathcal{F}_{i}$ as
$\mathcal{F}_{i}\doteq\operatorname{proj}_{\mathcal{Z}_{i}}\mathcal{X}_{i}.$ (15)
An example of the projection of a three-dimensional polytope onto the xy-plane is shown in Figure 2.
Figure 2: Example of the projection $\operatorname{proj}(\mathcal{X})$ of a three dimensional polytope $\mathcal{X}$ on the xyxy-plane.
III Complexity Reduction using Constrained Zonotopes
Different set representations offer different trade-offs in terms of representational accuracy, compactness, and suitability for certain operations, such as set projection, linear maps, and intersection. Therefore, the conversion into an appropriate set representation is critical to perform projection, i.e. aggregation, in high-dimensional problems such as multi-stage OPF.
III-A Approximation of Feasible Operating Region
Given an affine set representation of the multi-stage OPF, the feasible set $\mathcal{X}_{i}$ is a convex and bounded polytope in half-space representation (H-rep) and can be expressed as
$\mathcal{P}\doteq\biggl\{x\in\mathbb{R}^{n}\ \bigg|\ Ax\leq b\biggr\},$
with a constraint matrix $A\in\mathbb{R}^{m\times n}$ and a constraint vector $b\in\mathbb{R}^{m}$. This formulation enables the projection of the feasible set onto coupling variables using methods from computational geometry, such as Fourier–Motzkin elimination or vertex enumeration [21, pp. 2-3]. However, for multi-stage flexibility aggregation with storage units, the increased problem dimension renders these algorithms computationally intractable.
III-B Reformulation of the Feasible Set
Zonotopes are well suited for projections, also in high dimensions. They are defined as
$\mathcal{Z}\doteq\biggl\{c+\sum_{i=1}^{n_{g}}\alpha_{i}G_{(\cdot,i)}\ \bigg|\ \alpha_{i}\in[-1,1]\biggr\},$
with a center vector $c\in\mathbb{R}^{n}$, a generator matrix $G\in\mathbb{R}^{n\times n_{g}}$ and the scalar internal factors $\alpha_{i}$. Zonotope projection is performed via a simple sparse linear map and therefore can be computed very efficiently. In the following, the shorthand notation $\mathcal{Z}=\langle c,G\rangle_{\mathcal{Z}}$ is used for zonotopes. However, zonotopes are limited to model centrally symmetric convex polytopes. To address this limitation Constrained Zonotopes ($\mathcal{CZ}\mathrm{s}$) extend zonotopes by adding linear constraints to the internal factors $\alpha_{i}$ [16]. This results in the definition
$\mathcal{CZ}\doteq\biggl\{c+\sum_{i=1}^{n_{g}}\alpha_{i}G_{(\cdot,i)}\ \bigg|\ \sum_{i=1}^{n_{g}}\alpha_{i}A_{(\cdot,i)}=b,\quad\alpha_{i}\in[-1,1]\biggl\},$
with center vector $c\in\mathbb{R}^{n}$, generator matrix $G\in\mathbb{R}^{n\times n_{g}}$, constraint matrix $A\in\mathbb{R}^{m\times n_{g}}$, and constraint vector $b\in\mathbb{R}^{m}$. In the following, the shorthand notation $\mathcal{CZ}=\langle c,G,A,b\rangle_{\mathcal{CZ}}$ is used for constrained zonotopes. Without limitations on the number of generators $n_{g}$ and the number of constraints $m$, $\mathcal{CZ}\mathrm{s}$ can express arbitrary convex polytopes and offer many of the computational advantages of zonotopes, like projection via matrix multiplication [16].
The conversion of the feasible set $\mathcal{X}_{i}$ of the multi-stage OPF into a $\mathcal{CZ}$ can be decomposed into three sequential stages, where the first two steps can be considered as an offline computation to achieve the desired set representation as $\mathcal{CZ}$ whereas the third step is set projection operation itself.
Determine a bounding zonotope
Given a feasible set as convex and bounded polytope one can determine upper and lower bounds for each variable by solving two Linear Programms (LPs) for each variable $x_{i,k}$. This allows to determine the vectors of upper and lower variable bounds $ub_{x}$ and $lb_{x}$, which define the bounding zonotope $Z_{b}$ with center $c$ and generator matrix $G$ that over-approximates the given polytope $Z_{b}=\langle c,G\rangle_{Z}\subseteq\mathcal{P}$ as follows:
$c$ $=\tfrac{1}{2}(ub_{x}+lb_{x})$ (16a)
$G$ $=\text{diag}(\tfrac{1}{2}(ub_{x}-lb_{x}))$ (16b)
$Z_{b}$ $=\langle c,\ G\rangle.$ (16c)
However, as shown in Subsection V-B selecting the variable bounds sufficiently large is adequate without loss of accuracy for any $\tilde{lb}$ and $\tilde{ub}$ as long as
$\tilde{lb}_{x_{i},k}\leq lb_{x_{i},k}\leq x_{i,k}\leq ub_{x_{i},k}\leq\tilde{ub}_{x_{i},k}.$ (17)
While significantly reducing computation times in high dimensional problems such as multi-stage OPF, this may result in a highly over-approximative $Z_{b}$. However, the following step of the conversion into a $\mathcal{CZ}$ applies all constraints defining the feasible set $\mathcal{X}_{i}$ to the internal factors $\alpha$ of the $\mathcal{CZ}$ representation, i.e. effectively constraining the symmetric $Z_{b}$ to model the feasible set exactly.
Adding constraints to the bounding zonotope
Given the feasible set $\mathcal{X}_{i}$ of the multi-stage OPF as polytope $\mathcal{P}$ in H-rep, it is defined as an intersection of a finite number of halfspaces $H_{i}=\{x\in\mathbb{R}^{n}\mid h^{\top}x\leq\zeta\}$. To reconstruct the polytope as $\mathcal{CZ}$, the bounding zonotope is successively intersected with each halfspace $H_{i}$. Based on the method in [22], each intersection updates the $\mathcal{CZ}$ as follows:
$\mathcal{CZ}_{H}=\langle c,\begin{bmatrix}G&0\end{bmatrix},\tilde{A},\tilde{b}\rangle_{\mathcal{CZ}},$ (18)
with
$\tilde{A}=\begin{bmatrix}A&0\\ h^{\top}G&\frac{d_{m}}{2}\end{bmatrix},\quad\tilde{b}=\begin{bmatrix}b\\ \zeta-h^{\top}c-\frac{d_{m}}{2}\end{bmatrix}$, (19)
where $d_{m}$ is given by
$d_{m}=\zeta-h^{\top}c+\sum_{i=1}^{n_{g}}|h^{\top}g_{i}|.$ (20)
This step assumes that the provided polytope is irredundant, so that no feasibility checks are required before intersection. Otherwise, it is beneficial to keep the representation compact by solving an LP to check whether a halfspace intersects the current $\mathcal{CZ}$ and therefore has to be added to the representation [22]. The proposed approach is summarized in Algorithm 1.
Algorithm 1 Complexity reduction using $\mathcal{CZ}\mathrm{s}$.
Require Bounding zonotope $\mathcal{Z}=\langle c,G\rangle$, halfspaces $\mathcal{H}=\{H_{1},\dots,H_{m}\}$, defining polytope $P$
Ensure $\mathcal{CZ}=\langle c,G,A,b\rangle$
Initialize $\mathcal{CZ}=\langle c,G,A,b\rangle$ with empty $A,b$
Offline computation:
1:for all $H_{i}$ modeling static constraints (5)-(2), (4), (9) do
2:  Compute (20)
3:  Update $\mathcal{CZ}$ as described in (18) and (19)
4:end for
Online computation:
1:for all $H_{i}$ modeling dynamic constraints like (3) do
2:  Compute (20)
3:  Update $\mathcal{CZ}$ as described in (18) and (19)
4:end for
Return $\mathcal{CZ}$
The addition of a zero column to $G$ in (18) does not affect the result of $h^{\top}G$. This allows to parallelize the calculation of the rows for $A,b$ of the $\mathcal{CZ}$, which greatly benefits the required computation time.
Projection of the resulting constrained zonotope
Finally, once the $\mathcal{CZ}$ is constructed, projection onto the desired subspace is performed via the sparse linear map $M_{p}$, yielding:
$M_{p}\otimes\mathcal{CZ}=\langle M_{p}c,\ M_{p}G,\ A,\ b\rangle.$ (21)
This projection operation utilizes the computational benefits of the $\mathcal{CZ}$ representation, enabling fast and scalable dimensionality reduction via linear map.
IV Intersection with Additional Linear Constraints
For more dynamic use cases, it is essential to limit the computation cost required to consider changes in the feasible set $\mathcal{X}_{i}$ due to control inputs. Given the inherent separation of the set reformulation from the projection calculation in the presented method, most of the computation cost can be performed as an offline computation. Further, we can utilize that $\mathcal{CZ}\mathrm{s}$ allow to calculate halfspace and hyperplane intersections. This allows to add additional constraints to the $\mathcal{CZ}$ after initial setup, which is especially beneficial for increasing problem size, as most of the problem remains identical per time step. The static constraints (5)-(2), (4), (9) in $\mathcal{X}_{i}$ result in most of the offline computation time, which can be performed once. Then, the resulting $\mathcal{CZ}$ is modified as needed with reduced computation time, i.e. setpoint constraints can be adapted dynamically by adding the respective constraints via intersection each iteration. For instance, this enables the integration of revised forecast data in (3). The computational efficiency of these intersections is evaluated in Subsection V-B.
V Numerical Case Study
All numerical results in this section are performed on a system with an eight core and 16 thread Ryzen 9800X3D CPU paired with 64GB of 6000MHz RAM. Provided computation times are the average of ten executions. We use JuMP.jl [23] and MOSEK [24] to solve optimization problems. For convex sets, Polyhedra.jl allows to extract the set directly form the JuMP.jl model as polytope in H-representation. For the problem formulation, we consider LinDistFlow with and without loss linearization as as discussed in Section II. This allows to achieve a convex approximation while maintaining the crucial information of reactive power and voltage magnitude in the DSO grid model. As the underlying problem, we consider a multi-stage Optimal Power Flow (MS-OPF) problem, which allows flexibililty aggregation across multiple time steps and serves to evaluate computational scalability for horizons of up to $N=96$.
V-A Comparison of Computation Times
First, we consider a small-scale, 4-bus radial distribution grid adapted from MATPOWER, called case4dist [25]. The grid has been extended to include two additional renewable generators at buses three and four, thereby increasing the system’s overall active and reactive power flexibility. This small test case allows the comparison of the presented method with the baseline approach of polytope projection, which can still be evaluated over a short varying time horizon of the MS-OPF. While this projection-based approach does not scale well to larger grids, it remains suitable for evaluating performance in this compact setting.
TABLE I: Computation times of the presented method and polytope projection on the 4-bus distribution grid model for up to $N=4$ timesteps. Fastest computation times in green.
Horizon NN Approach Offline [s] Online [s] Total [s]
1 $\mathcal{CZ}$ 0.000 657 250.000\,657\,25 $1.429\,99\text{\times}{10}^{-6}$ 0.000 658 650.000\,658\,65
[…]
2 $\mathcal{CZ}$ 0.002 958 739 90.002\,958\,739\,9 $7.8000\text{\times}{10}^{-7}$ 0.002 959 520.002\,959\,52
[…]
3 $\mathcal{CZ}$ 0.007 280 230.007\,280\,23 $1.02\text{\times}{10}^{-6}$ 0.007 281 250.007\,281\,25
[…]
4 $\mathcal{CZ}$ 0.011 396 410 000 000 0010.011\,396\,410\,000\,000\,001 $29\,999\,999\,999\,999\,9\text{\times}{10}^{-6} 0.011$ 397 440.011\,397\,44
[…]
Table I shows the computation times for the multi-stage OPF on the 4-bus distribution grid model for up to $N=4$ timesteps. The results highlight that for small problem sizes, the presented method achieves significantly faster computation times, especially for the online computation of the set projection onto the coupling variables of active and reactive power at point of interconnection. While both methods model the convex feasible set of the multi-stage OPF based on the LinDistFlow formulation with loss linearization, the presented method achieves faster total computation times for all time horizon lengths depicted in Table I as the polytope projection, where no offline computation time is needed. This highlights, that for small problem sizes even the set conversion itself can be performed online, as its computation time is sufficiently fast.
V-B Evaluation on Medium-sized Feeder
[…]
123456789101112131415
1234791211101456138
Figure 3: Topology of the 15-bus radial distribution grid with battery highlighted in green.
Approximation quality of reformulation based approaches
First, we demonstrate the equivalence of the reformulation of the feasible set as $\mathcal{CZ}$. Figure 4 shows that approximating the variable bounds does not impact the accuracy of the resulting $\mathcal{CZ}$. The FOR modeled as $\mathcal{CZ}$ $\mathcal{F}_{\mathcal{CZ}}^{LDF-LL}$ is equivalent to the respective FOR modeled as polytope $\mathcal{F}^{LDF-LL}$. As described in Subsection III-B, this holds as long as the bounding zonotope over-approximates the polytope and all constraints defining the polytope are added to the $\mathcal{CZ}$ representation in the next step of the conversion. Hence, all following computations use sufficiently large variable bounds to decrease computation time.
Figure 4: FOR at the slack bus modeled as polytope and as $\mathcal{CZ}$ using sufficiently large variable bounds.
TABLE II: Computation times of the presented method on the 15-bus distribution grid model for up to $N=96$ timesteps.
Horizon NN Offline [s] Online [s] Total [s]
12 1.178 872 261.178\,872\,26 $2.952\,000\,000\,000\,000\,2\text{\times}{10}^{-5}$ 1.178 901 781.178\,901\,78
24 5.488 087 765.488\,087\,76 0.000 113 950.000\,113\,95 5.488 201 715.488\,201\,71
36 23.795 031 979 999 99723.795\,031\,979\,999\,997 0.000 110 689 999 999 999 990.000\,110\,689\,999\,999\,999\,99 23.795 142 6723.795\,142\,67
48 61.923 130 689 999 99461.923\,130\,689\,999\,994 0.000 145 369 999 999 999 970.000\,145\,369\,999\,999\,999\,97 61.923 276 0661.923\,276\,06
60 124.313 936 030 000 01124.313\,936\,030\,000\,01 0.001 504 080.001\,504\,08 124.315 440 11124.315\,440\,11
72 267.925 450 57267.925\,450\,57 0.003 232 20.003\,232\,2 267.928 682 77267.928\,682\,77
84 311.894 214 320 000 06311.894\,214\,320\,000\,06 0.002 225 189 999 999 999 30.002\,225\,189\,999\,999\,999\,3 311.896 439 51311.896\,439\,51
96 535.231 617 39535.231\,617\,39 0.002 847 740.002\,847\,74 535.234 465 13535.234\,465\,13
Table II shows the respective timeseries of required computation times for the reformulation of the feasible set as $\mathcal{CZ}$. For the given results, the constraint intersection is performed in parallel. The results clearly show that despite the fact that the computation time scales noticeably, it remains suitable for offline computation on the used test problem, even for a considered time horizon length of $N=96$. Further, $\mathcal{CZ}$ projection can be performed in sub-second times on the 15-bus distribution grid case, even for large time horizons $N$. This highlight the suitability for the presented method to reduce complexity in ADP. Paired with acceptable conversion computation times, this allows the application of ADP to considerably larger problem sizes, by using the presented method to model the feasible sets of the sub problems as $\mathcal{CZ}$.
Computation times for set modification
To reduce the repeated necessity of performing the set conversion due to small changes in the feasible set, e.g. set point alterations, this section briefly evaluates the modification of an existing $\mathcal{CZ}$ with additional halfspace or hyperplane intersection as discussed in Section IV.
TABLE III: Computation times required to add an additional constraint to an existing $\mathcal{CZ}$ for up to $N=96$ timesteps.
Horizon NN Computation time [s]
12 0.033 474 930.033\,474\,93
24 0.245 869 019 999 999 940.245\,869\,019\,999\,999\,94
36 0.256 886 150 000 000 040.256\,886\,150\,000\,000\,04
48 0.442 441 160.442\,441\,16
60 1.380 779 31.380\,779\,3
72 3.699 329 563.699\,329\,56
84 4.824 102 814.824\,102\,81
96 6.559 998 849 999 999 56.559\,998\,849\,999\,999\,5
Table III depicts the time necessary to add an individual affine constraint to a $\mathcal{CZ}$ modeling the multi-stage OPF for increasing time horizon lengths $N$. The results show, that adding additional affine constraints is in fact viable as the required computation time remains sufficiently low, i.e. far below the considered time step size of fifteen minutes, even for large problem size.
These result highlight the potential use of these set representations for dynamic sets. For instance, this can be used in ADP to model a feasible set for a fixed set of constraints and add more dynamic constraints, e.g. time-varying generator setpoints, via further intersection as necessary.
Time-dependent feasible operating region
Given the significantly reduced computation times, we can now consider a MS-OPF problem on the 15-bus distribution grid case and therefore the FOR at point of interconnection for the coupling variables of active and reactive power along the time horizon. Figure 5 shows the available active and reactive power values of the multidimensional FOR at time step $k=2$, denoted by $p_{1,2}(2)$ and $q_{1,2}(2)$, depending on the provided active power at the point of interconnection at $k=1$, $p_{1,2}(1)$. One can see that the $p_{1,2}(2)$$q_{1,2}(2)$ area becomes significantly smaller with increasing $p_{1,2}(1)$. This is due to the battery dynamics (12), i.e. the available energy in the battery is lowered for timestep $k=2$ depending on the power drawn in timestep $k=1$. Otherwise, the available flexibility remains consistent as the loads as well as generation units in the 15-bus distribution grid model are considered to be static over time.
Figure 5: Dependency of $p_{1,2}(1)$ FOR of active and reactive power at the following timestep.
This highlights the necessity to consider the multi-stage OPF for the DSO subproblems in order to ensure a feasible solution for consecutive timesteps and to improve flexibilty utilization.
VI Conclusion and Future Work
This paper addresses the central computational bottleneck arising in ADP, which is the efficient computation of set projections. By utilizing $\mathcal{CZ}\mathrm{s}$ as beneficial set representations we can significantly reduce the computation times required for flexibility aggregation via set projection of convex feasible sets. This allows to perform time-dependent flexibility aggregation in the setting of multi-stage OPF and on large-scale problems without loss of accuracy. Further we show that most of the computation can be performed offline and set modifications can be considered efficiently.
Future work will focus on further application in the proposed use cases and the extension of the presented method to non-convex OPF formulations.
[…]